OpenStudy (anonymous):

The equation of line CD is (y−3) = − 2 (x − 4). What is the slope of a line perpendicular to line CD? A.1 over 2 B. 2 C. negative 1 over 2 D. −2

OpenStudy (anonymous):

you know what the slope of this line is ?

OpenStudy (anonymous):

hint: it is in the "point-slope" form, so you can read the slope directly from the equation

OpenStudy (haseeb96):

when you will solve the above equation you will get y-3 =-2x + 8 y = -2x -11.....eq#1 here -2 is your slope now for perpendicular slope will be +1/2 now put 1/2 at the place of -2

OpenStudy (haseeb96):

y= 1/2 x -11 2y = x - 22 x-2y-22=0 is the correct answer

OpenStudy (anonymous):

thats not one of the answer choices??????

OpenStudy (haseeb96):

A is the correct answer I just take out the equation of perpendicular line

OpenStudy (anonymous):

i chose that one last time and it said i was wrong

OpenStudy (haseeb96):



OpenStudy (haseeb96):

A is correct answer I am 100% sure

OpenStudy (haseeb96):

how -1/2 - signs will be cancel from each sides

OpenStudy (anonymous):

you're correct! option A (1/2)